OpenStudy (anonymous):

Find the area of a rhombus with diagonals measuring 8.6 in and 15.3 in. A. 65.79 in2 B. 131.58 in2 C. 32.90 in2 D. 23.9 in2

OpenStudy (anonymous):

Area= 1/2 times d1 times d2 where d1 and d2 are the diagonals

Directrix (directrix):

The area of a rhombus can be found by taking 1/2 the product of its diagonal lengths. A = 1/2 * 8.6 * 15.3 A = ?

OpenStudy (anonymous):

131.58 @Directrix

Directrix (directrix):

I did not get that. --> I think you forgot to take 1/2 of 131.58. See attachment for what the Google calculator cranked out.

Directrix (directrix):

A = 1/2 * 8.6 * 15.3 = 1/2 * 131.58 = ?